OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. a19 = -58, a21 = -164

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

hey:=) we need common difference and first term to write an equation arithmetic equation \[\huge\rm a_n = a_1 + (n-1)d\] where a_1 = first term d=common difference n= term that we have to find

Nnesha (nnesha):

given terms are a_{19}= -58 a_{21} = -164 plug this into the equation \[\large\rm a_\color{ReD}{{19}}=a_1 +(\color{ReD}{19}-1)d\] 19th term so we substitute n for 19 and a_{19} equal to -58 we can replace a_{19} with -58 \[\large\rm -58=a_1 +(\color{ReD}{19}-1)d\] same for a_{21} write an equation when n =21

Nnesha (nnesha):

your turn:=) write an equation when n =21 just like i did for n=19 :=)

Nnesha (nnesha):

make sense ??

OpenStudy (anonymous):

lets see, -164 = a1 (21 - 1)d ?

Nnesha (nnesha):

perfect! \[\huge\rm -164 = a_1 +(\color{Red}{21-1})d\] solve parentheses \[\large\rm -164 = a_1 +(\color{Red}{20})d\] this is our 2nd equation first one is \[\huge\rm -58 = a_1 +(\color{Red}{19-1})d\] \[\large\rm -58 = a_1 +(\color{Red}{18})d\]

Nnesha (nnesha):

\[\large\rm -58 = a_1 + \color{Red}{18} d\]\[\large\rm -164 = a_1 + \color{Red}{20} d\] we will use elimination method to solve this equation :=) so subtract equation one from equation 2